# The Bohlin variant of the Eisenhart lift

### v1

@@
Tomsk State University of Control Systems and Radioelectronics,
Lenin ave. 40, 634050 Tomsk, Russia
@@ 1. Introduction
$L(x,\dot{x})dt^{2}+dtds,$
where $L(x,\dot{x})$ is the Lagrangian density, $t$ is the temporal coordinate, and $s$ is an extra variable [1, 2]. The latter is needed in order to reproduce the conventional nonrelativistic energy conservation condition within the relativistic framework. At the level of equations of motion, $s$ links to the action variable sometimes adopted in classical mechanics. Thus, given a dynamical system with $d$ degrees of freedom $x_{i}$, $i=1,\dots,d$, the resulting space–time is (d+2)(d+2)–dimensional and the metric is of the Lorentzian signature.
@@
If a Lagrangian density explicitly depends on time, $L(t,x,\dot{x})$, the metric is referred to as the Brinkmann–type metric or a plane–fronted wave with parallel rays (the pppp–wave).
@@
The importance of the Killing vectors and tensors is hard to overestimate. They give a clue to establishing complete integrability of the geodesic equations formulated in various black hole spacetimes, as well as allow one to separate variables in the Hamilton–Jacobi, Klein–Gordon and Dirac equations in strong gravitational fields (for a review and further references see [3]).
Apart from the aesthetic appeal, the Eisenhart lift provides an efficient means of constructing novel space–times with hidden symmetries. By applying the method to various integrable models one can build a plethora of examples of such a kind [4][16]. Some of them provide exact solutions to the Einstein equations [4, 11, 12, 15]. Worth mentioning also is a considerable body of recent work on application of the Eisenhart lift to cosmology (see e.g. [17][20] and references therein).
The goal of this work is to study a variant of the Eisenhart lift, which is inspired by the Bohlin transformation [21] (for a review see [22]) connecting the planar harmonic oscillator to the Kepler problem. Given a dynamical system with $d$ degrees of freedom, a simple conformally flat metric on a (d+2)(d+2)–dimensional space–time of the Lorentzian signature is proposed, timelike geodesics of which include the original Newton equation.
@@ 2. The Eisenhart lift
Consider a conservative dynamical system with $d$ degrees of freedom $x_{i}$, $i=1,\dots,d$, and let $t$ be the temporal variable. Assuming that potential energy $V(x)$ can be factorized $V(x)=gU(x)$, where $g$ is a constant having the dimension of energy and $U(x)$ is a positive dimensionless function, [^1] Because the coordinates $x_{i}$ have the dimension of length and $U(x)$ is assumed to be dimensionless, the latter will necessarily involve a (coupling) constant with the dimension of length. By this reason, a numerical value of $g$ entering the original potential $V(x)=gU(x)$ is of no great importance. the Newton equation
@@
Finally, assuming the original dynamical system (1) admits a polynomial integral of motion
@@
one can uplift it to a Killing tensor [^3] For a discussion of conformal Killing tensors within the Eisenhart framework see e.g. [7, 10]. of rank $n$ just by multiplying $\mathcal{I}$ with ${\left(\frac{dt}{d\lambda}\right)}^{n}$ and representing the latter in the form
$K_{A_{1}\dots A_{n}}(y)\frac{dy^{A_{1}}}{d\lambda}\dots\frac{dy^{A_{n}}}{d\lambda}.$
The resulting expression will automatically obey $\nabla_{(A_{1}}K_{A_{2}\dots A_{n+1})}=0$.
@@ 3. The Bohlin variant of the Eisenhart lift
Note that, if (10) happens to be a polynomial integral of motion of a dynamical system at hand, one can turn it into a Killing tensor of the Bohlin metric. The uplift is less straightforward than within the Eisenhart framework, however. Because Newton’s equation (1) holds invariant under the time reversal transformation $t\to-t$, the corresponding first integrals involve monomials in momenta which are all of even/odd power. Substituting $p_{i}=\frac{dx_{i}}{dt}=\frac{U}{\alpha}\frac{dx_{i}}{d\tau}$ into the integrals of motion and multiplying, where necessary, with the unity $1=\frac{1}{c^{2}}g_{AB}\frac{dy^{A}}{d\tau}\frac{dy^{B}}{d\tau}$, one obtains a homogeneous function $K_{A_{1}\dots A_{n}}\frac{dy^{A_{1}}}{d\tau}\dots\frac{dy^{A_{n}}}{d\tau}$, from which components of the Killing tensor are read off.
@@ 4. Examples
where $a$ is a dimensionless constant and $b_{i}$ are constants having the dimension of inverse length. When Bohlin uplifted, it results in a metric solving the vacuum Einstein equations in the presence of a cosmological constant [^5] As a matter of fact, (23) yields a metric defined on a patch of the (d+2)(d+2)–dimensional anti de Sitter space, because $R_{ABCD}=(b_{i}b_{i})\left(g_{AC}g_{BD}-g_{AD}g_{BC}\right)$.
@@
$V(x_{1},\dots,x_{d})=\frac{1}{2}\sum_{i\neq j}^{d}\frac{g^{2}}{{(x_{i}-x_{j})}^{2}},$ (24)
@@
$I_{1}=\sum_{i=1}^{d}p_{i},\quad I_{2}=\frac{1}{2}\sum_{i=1}^{d}p_{i}^{2}+\sum_{i<j}^{d}\frac{g^{2}}{{(x_{i}-x_{j})}^{2}},\quad I_{3}=\frac{1}{3!}\left(\sum_{i=1}^{d}p_{i}^{3}+3g^{2}\sum_{i<j}^{d}\frac{p_{i}+p_{j}}{{(x_{i}-x_{j})}^{2}}\right),$
$I_{4}=\frac{1}{4!}\left(\sum_{i=1}^{d}p_{i}^{4}+4g^{2}\sum_{i<j}^{d}\frac{p_{i}^{2}+p_{j}^{2}+p_{i}p_{j}}{{(x_{i}-x_{j})}^{2}}+2g^{4}\sum_{i<j}^{d}\frac{1}{{(x_{i}-x_{j})}^{4}}+\right$.
$\qquad\qquad\left.+4g^{4}\sum_{i\neq j,i\neq k,j<k}^{d}\frac{1}{{(x_{i}-x_{j})}^{2}{(x_{i}-x_{k})}^{2}}\right).$ (25)
@@
$U(x_{1},\dots,x_{d})=\frac{1}{2}\sum_{i\neq j}^{4}\frac{L^{2}}{{(x_{i}-x_{j})}^{2}},$
@@
$K_{tsi}=\frac{1}{2}L^{2}U^{2}\sum_{a,b,a\neq b}^{4}\frac{1}{{(x_{a}-x_{b})}^{2}},$
$K_{ijp}=U^{3}\delta_{ip}\delta_{jp}-\frac{1}{2}L^{2}U^{2}\left(\delta_{ij}\sum_{a\neq p}^{4}\frac{1}{{(x_{a}-x_{p})}^{2}}+\delta_{jp}\sum_{a\neq i}^{4}\frac{1}{{(x_{a}-x_{i})}^{2}}+\delta_{pi}\sum_{a\neq j}^{4}\frac{1}{{(x_{a}-x_{j})}^{2}}\right),$
@@
$K_{ttss}=\frac{1}{6}L^{4}U^{2}\left(\sum_{a,b,a\neq b}^{4}\frac{1}{{(x_{a}-x_{b})}^{4}}+2\sum_{a,b,c,a\neq b\neq c}^{4}\frac{1}{{(x_{a}-x_{b})}^{2}{(x_{a}-x_{c})}^{2}}\right),$
$K_{tsij}=-\frac{1}{12}L^{4}U^{2}\delta_{ij}\left(\sum_{a,b,a\neq b}^{4}\frac{1}{{(x_{a}-x_{b})}^{4}}+2\sum_{a,b,c,a\neq b\neq c}^{4}\frac{1}{{(x_{a}-x_{b})}^{2}{(x_{a}-x_{c})}^{2}}\right)$
$+\frac{1}{6}L^{2}U^{3}\sum_{a,b,a\neq b}^{4}\frac{\left(2\delta_{ai}\delta_{aj}+\delta_{ai}\delta_{bj}\right)}{{(x_{a}-x_{b})}^{2}},$
@@
$+\frac{1}{12}L^{4}U^{2}\left(\delta_{ij}\delta_{kp}+\delta_{ik}\delta_{pj}+\delta_{ip}\delta_{jk}\right)\left(\sum_{a,b,a\neq b}^{4}\frac{1}{{(x_{a}-x_{b})}^{4}}+2\sum_{a,b,c,a\neq b\neq c}^{4}\frac{1}{{(x_{a}-x_{b})}^{2}{(x_{a}-x_{c})}^{2}}\right)$
$-\frac{1}{6}L^{2}U^{3}\left(2\sum_{a,b,a\neq b}^{4}\frac{\left(\delta_{ai}\delta_{aj}\delta_{kp}+\delta_{ai}\delta_{ak}\delta_{pj}+\delta_{ai}\delta_{ap}\delta_{jk}+\delta_{aj}\delta_{ak}\delta_{pi}+\delta_{aj}\delta_{ap}\delta_{ik}+\delta_{ak}\delta_{ap}\delta_{ij}\right)}{{(x_{a}-x_{b})}^{2}}\right$.
$\left.+\sum_{a,b,a\neq b}^{4}\frac{\left(\delta_{ai}\delta_{bj}\delta_{kp}+\delta_{ai}\delta_{bk}\delta_{pj}+\delta_{ai}\delta_{bp}\delta_{jk}+\delta_{aj}\delta_{bk}\delta_{pi}+\delta_{aj}\delta_{bp}\delta_{ik}+\delta_{ak}\delta_{bp}\delta_{ij}\right)}{{(x_{a}-x_{b})}^{2}}\right).$
@@ 5. Conclusion
To summarize, inspired by the Bohlin transformation relating the planar harmonic oscillator to the Kepler problem, a variant of the Eisenhart lift was proposed, in which a Lagrangian conservative dynamical system with $d$ degrees of freedom was embedded into timelike geodesics of a conformally flat metric defined on a (d+2)(d+2)–dimensional space–time of the Lorentzian signature. The (anti) de Sitter space–time fitted nicely into the framework. The uplift was used to construct novel conformally flat metrics admitting higher rank Killing tensors.
Let us recapitulate the main features of the construction. A potential of the original dynamical system determines the conformal factor of the metric, as well as it relates the coordinate time to the proper time. Apart from Cartesian coordinates and a temporal variable, an extra spatial coordinate is needed in order to reproduce the conventional nonrelativistic energy conservation condition within the relativistic framework. In general, the metric admits three Killing vector fields, which form Lie algebra of the Poincaré group in $1+1$ dimensions. None of them is covariantly constant and, hence, the space–time does not belong to the Kundt class. If a dynamical system at hand admits an integral of motion polynomial in momenta, it can be uplifted to a Killing tensor. The uplift is different from that adopted within the conventional Eisenhart framework and essentially relies upon the condition that the geodesic is timelike.
@@ Acknowledgements
This work was supported by the RF Ministry of Science and Higher Education under the project FSWW-2026-0010.

### v2

@@ 1. Introduction
$\mathcal{L}(x,\dot{x})dt^{2}+dtds,$
where $\mathcal{L}(x,\dot{x})$ is the Lagrangian density, $t$ is the temporal coordinate, and $s$ is an extra variable [1, 2]. The latter is needed in order to reproduce the conventional nonrelativistic energy conservation condition within the general relativistic framework. At the level of equations of motion, $s$ links to the action variable sometimes adopted in classical mechanics. Thus, given a dynamical system with $d$ degrees of freedom $x_{i}$, $i=1,\dots,d$, the resulting space–time is (d+2)(d+2)–dimensional and the metric is of the Lorentzian signature.
@@
If a Lagrangian density explicitly depends on time, $\mathcal{L}(t,x,\dot{x})$, the metric is referred to as the Brinkmann–type metric or a plane–fronted wave with parallel rays (the pppp–wave).
@@
The importance of the Killing vectors and tensors is hard to overestimate. They give a clue to establishing complete integrability of the geodesic equations formulated in various black hole space–times, as well as allow one to separate variables in the Hamilton–Jacobi, Klein–Gordon and Dirac equations in strong gravitational fields (for a review and further references see [3]).
Apart from the aesthetic appeal, the Eisenhart lift provides an efficient means of constructing novel space–times with hidden symmetries. By applying the method to various integrable models, one can build a plethora of examples of such a kind [4][16]. Some of them provide exact solutions to the Einstein equations [4, 11, 12, 15]. Worth mentioning also is a considerable body of recent work on application of the Eisenhart lift to cosmology (see e.g. [17][20] and references therein).
The goal of this paper is to study a variant of the Eisenhart lift, which is inspired by the Bohlin transformation [21] (for a review see [22]) connecting the planar harmonic oscillator to the Kepler problem. Given a dynamical system with $d$ degrees of freedom, a simple conformally flat metric on a (d+2)(d+2)–dimensional space–time of the Lorentzian signature is proposed, timelike geodesics of which include the original Newton equation.
@@ 2. The Eisenhart lift
Consider a conservative dynamical system with $d$ degrees of freedom $x_{i}$, $i=1,\dots,d$, and let $t$ be the temporal variable. Assuming that potential energy $V(x)$ can be factorized $V(x)=gU(x)$, where $g$ is a constant having the dimension of energy and $U(x)$ is a positive dimensionless function, [^1] Because the coordinates $x_{i}$ have the dimension of length and $U(x)$ is assumed to be dimensionless, the latter will necessarily involve a (coupling) constant with the dimension of length. By this reason, a numerical value of $g$ entering the original potential $V(x)=gU(x)$ is of no great importance. In what follows, we refer to $U$ as the prepotential. the Newton equation
@@
Finally, assuming the original dynamical system (1) admits a polynomial integral of motion of degree $n$ in velocities
@@
one can uplift it to a Killing tensor [^3] For a discussion of conformal Killing tensors within the Eisenhart framework see e.g. [7, 10]. of rank $n$. Recall that the geodesic equations (2), (5) imply that the coordinate time $t$ and the affine parameter $\lambda$ are affinely related $\frac{dt}{d\lambda}=\alpha$. By multiplying the conserved charge $\mathcal{I}$ with the constant $\alpha^{n}={\left(\frac{dt}{d\lambda}\right)}^{n}$, where $n$ is the highest power of the velocity in (10), one can bring the resulting expression to the homogeneous form
$K_{A_{1}\dots A_{n}}(y)\frac{dy^{A_{1}}}{d\lambda}\dots\frac{dy^{A_{n}}}{d\lambda},$
where $y^{A}=(ct,s,x_{1},\dots,x_{d})$, from which the components of the Killing tensor obeying $\nabla_{(A_{1}}K_{A_{2}\dots A_{n+1})}=0$ are read off.
@@ 3. The Bohlin variant of the Eisenhart lift
Note that, if (10) happens to be a polynomial integral of motion of a dynamical system at hand, one can turn it into a Killing tensor of the Bohlin metric. The uplift is less straightforward than within the Eisenhart framework, however. Because Newton’s equation (1) holds invariant under the time reversal transformation $t\to-t$, the corresponding first integrals involve monomials in momenta which are all of even/odd power. Substituting $p_{i}=\frac{dx_{i}}{dt}=\frac{U}{\alpha}\frac{dx_{i}}{d\tau}$ into an integral of motion of degree $n$ in momenta and multiplying, where necessary, with the unity $1={\left(\frac{1}{c^{2}}g_{AB}\frac{dy^{A}}{d\tau}\frac{dy^{B}}{d\tau}\right)}^{m}$, where $m$ is a natural number, one can turn the latter into a homogeneous function $K_{A_{1}\dots A_{n}}\frac{dy^{A_{1}}}{d\tau}\dots\frac{dy^{A_{n}}}{d\tau}$ of degree $n$ in the velocities $\frac{dy^{A}}{d\tau}$, from which components of the Killing tensor $K_{A_{1}\dots A_{n}}$ are read off. In contrast to the conventional Eisenhart lift, higher rank Killing tensors of the Bohlin–type metric will always involve tsts–components.
@@ 4. Examples
where $a$ is a dimensionless constant and $b_{i}$ are constants having the dimension of inverse length. When Bohlin uplifted, it results in a metric solving the vacuum Einstein equations in the presence of a cosmological constant [^5] As a matter of fact, (23) yields a metric defined on a patch of the (d+2)(d+2)–dimensional anti de Sitter space–time, because $R_{ABCD}=(b_{i}b_{i})\left(g_{AC}g_{BD}-g_{AD}g_{BC}\right)$. Anti de Sitter and de Sitter space–times thus fit nicely into the framework.
@@
$V(x_{1},\dots,x_{d})=\frac{1}{2}\sum_{i,j=1,i\neq j}^{d}\frac{g^{2}}{{(x_{i}-x_{j})}^{2}},$ (24)
@@
$I_{1}=\sum_{i=1}^{d}p_{i},\qquad I_{2}=\frac{1}{2}\sum_{i=1}^{d}p_{i}^{2}+\sum_{i,j=1,i<j}^{d}\frac{g^{2}}{{(x_{i}-x_{j})}^{2}},$
$I_{3}=\frac{1}{3!}\left(\sum_{i=1}^{d}p_{i}^{3}+3g^{2}\sum_{i,j=1,i<j}^{d}\frac{p_{i}+p_{j}}{{(x_{i}-x_{j})}^{2}}\right),$
$I_{4}=\frac{1}{4!}\left(\sum_{i=1}^{d}p_{i}^{4}+4g^{2}\sum_{i,j=1,i<j}^{d}\frac{p_{i}^{2}+p_{j}^{2}+p_{i}p_{j}}{{(x_{i}-x_{j})}^{2}}+2g^{4}\sum_{i,j=1,i<j}^{d}\frac{1}{{(x_{i}-x_{j})}^{4}}+\right$.
$\qquad\qquad\left.+4g^{4}\sum_{i,j,k=1,i\neq j,i\neq k,j<k}^{d}\frac{1}{{(x_{i}-x_{j})}^{2}{(x_{i}-x_{k})}^{2}}\right).$ (25)
@@
$U(x_{1},\dots,x_{4})=\frac{1}{2}\sum_{i,j,i\neq j}^{4}\frac{L^{2}}{{(x_{i}-x_{j})}^{2}},$
@@
$K_{tsi}=\frac{1}{2}L^{2}U^{2}\sum_{a=1,a\neq i}^{4}\frac{1}{{(x_{a}-x_{i})}^{2}},$
$K_{ijp}=U^{3}\delta_{ip}\delta_{jp}-\frac{1}{2}L^{2}U^{2}\left(\delta_{ij}\sum_{a=1,a\neq p}^{4}\frac{1}{{(x_{a}-x_{p})}^{2}}+\delta_{jp}\sum_{a=1,a\neq i}^{4}\frac{1}{{(x_{a}-x_{i})}^{2}}\right$.
$\left.\qquad\quad+\delta_{pi}\sum_{a=1,a\neq j}^{4}\frac{1}{{(x_{a}-x_{j})}^{2}}\right),$
@@
$K_{ttss}=\frac{1}{6}L^{4}U^{2}\left(\sum_{a,b=1,a\neq b}^{4}\frac{1}{{(x_{a}-x_{b})}^{4}}+2\sum_{a,b,c=1,a\neq b\neq c}^{4}\frac{1}{{(x_{a}-x_{b})}^{2}{(x_{a}-x_{c})}^{2}}\right),$
$K_{tsij}=-\frac{1}{12}L^{4}U^{2}\delta_{ij}\left(\sum_{a,b=1,a\neq b}^{4}\frac{1}{{(x_{a}-x_{b})}^{4}}+2\sum_{a,b,c=1,a\neq b\neq c}^{4}\frac{1}{{(x_{a}-x_{b})}^{2}{(x_{a}-x_{c})}^{2}}\right)$
$\qquad\qquad+\frac{1}{6}L^{2}U^{3}\sum_{a,b=1,a\neq b}^{4}\frac{\left(2\delta_{ai}\delta_{aj}+\delta_{ai}\delta_{bj}\right)}{{(x_{a}-x_{b})}^{2}},$
@@
$+\frac{1}{12}L^{4}U^{2}\left(\delta_{ij}\delta_{kp}+\delta_{ik}\delta_{pj}+\delta_{ip}\delta_{jk}\right)\left(\sum_{a,b=1,a\neq b}^{4}\frac{1}{{(x_{a}-x_{b})}^{4}}+2\sum_{a,b,c=1,a\neq b\neq c}^{4}\frac{1}{{(x_{a}-x_{b})}^{2}{(x_{a}-x_{c})}^{2}}\right)$
$-\frac{1}{6}L^{2}U^{3}\left(2\sum_{a,b=1,a\neq b}^{4}\frac{\left(\delta_{ai}\delta_{aj}\delta_{kp}+\delta_{ai}\delta_{ak}\delta_{pj}+\delta_{ai}\delta_{ap}\delta_{jk}+\delta_{aj}\delta_{ak}\delta_{pi}+\delta_{aj}\delta_{ap}\delta_{ik}+\delta_{ak}\delta_{ap}\delta_{ij}\right)}{{(x_{a}-x_{b})}^{2}}\right$.
$\left.+\sum_{a,b=1,a\neq b}^{4}\frac{\left(\delta_{ai}\delta_{bj}\delta_{kp}+\delta_{ai}\delta_{bk}\delta_{pj}+\delta_{ai}\delta_{bp}\delta_{jk}+\delta_{aj}\delta_{bk}\delta_{pi}+\delta_{aj}\delta_{bp}\delta_{ik}+\delta_{ak}\delta_{bp}\delta_{ij}\right)}{{(x_{a}-x_{b})}^{2}}\right).$
It is straightforward to verify that the Killing tensors are irreducible, i.e. they do not decompose into a combination of the Killing vector fields (13) and the metric (12).
@@ 5. Conclusion
To summarize, inspired by the Bohlin transformation relating the planar harmonic oscillator to the Kepler problem, a variant of the Eisenhart lift was proposed, in which a Lagrangian conservative dynamical system with $d$ degrees of freedom was embedded into timelike geodesics of a conformally flat stationary metric defined on a (d+2)(d+2)–dimensional space–time of the Lorentzian signature. The (anti) de Sitter space–time fitted nicely into the framework. The uplift was used to construct novel conformally flat metrics admitting higher rank Killing tensors.
Let us recapitulate the main features of the construction. A potential of the original dynamical system determines the conformal factor of the metric, as well as it relates the coordinate time to the proper time. Apart from Cartesian coordinates and a temporal variable, an extra spatial coordinate is needed in order to reproduce the conventional nonrelativistic energy conservation condition within the general relativistic framework. In general, the metric admits three Killing vector fields, which form the Lie algebra of the Poincaré group in $1+1$ dimensions. None of them is covariantly constant and, hence, the space–time does not belong to the Kundt class. If a dynamical system at hand admits an integral of motion polynomial in momenta, it can be uplifted to a Killing tensor. The uplift is different from that adopted within the conventional Eisenhart framework and essentially relies upon the condition that the geodesic is timelike.
@@ Acknowledgements
This work was supported by the RF Ministry of Science and Higher Education under the project FSWW-2026-0046.